# Deep Proxy Causal Learning and its Application to Confounded Bandit Policy Evaluation

### v4

@@ 3.4 DFPV for Policy Evaluation
PCL methods can estimate not only the structural function, but also other causal quantities, such as Conditional Average Treatment Effect (CATE) [Singh, 2020] and Average Treatment Effect on Treated (ATT) [Deaner, 2018, Singh, 2020]. Beyond the causal context, [Tennenholtz et al., 2020] uses a PCL method to conduct the off-policy evaluation in a partially observable Markov decision process with discrete action-state spaces. Here, we present the application of PCL methods to bandit off-policy evaluation with continuous actions, and propose a method to leverage adaptive feature maps in this context.
@@ 4 Experiments
In this section, we report the empirical performance of the DFPV method. First, we present the results of estimating structural functions; we design two experimental settings for low-dimensional treatments and high-dimensional treatments, respectively. Then, we show the result of applying PCL methods to the bandit off-policy evaluation problem with confounding. We include the results for problems considered in prior work in Appendix E. The experiments are implemented using PyTorch [Paszke et al., 2019]. The code is included in the supplemental material. All experiments can be run in a few minutes on Intel(R) Xeon(R) CPU E5-2698 v4 2.20GHz.
@@
Our second structural function estimation experiment considers high-dimensional treatment variables. We test this using the dSprite dataset [Matthey et al., 2017], which is an image dataset described by five latent parameters (shape, scale, rotation, posX and posY). The images are $64\times 64=4096$-dimensional. Based on this, Xu et al. [2021] introduced the causal experiment, where the treatment is each figure, and the confounder is posY. Inspired by this, we consider the PCL setting that learns the same structural functions with nonlinear confounding, which is not possible to handle in the instrumental variable setting. Specifically, the structural function $f_{\mathrm{struct}}$ and outcome $Y$ are defined as
$f_{\mathrm{struct}}(a)=\frac{(\mathrm{vec}(\bm{B})^{\top}a)^{2}-3000}{500},\quad Y=\frac{1}{12}(\mathtt{posY}-0.5)f_{\mathrm{struct}}(A)+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,0.5),$
@@
We compare the DFPV method to three competing methods, namely KPV [Mastouri et al., 2021], PMMR [Mastouri et al., 2021], and an autoencoder approach derived from the CEVAE method [Louizos et al., 2017]. In KPV, the bridge function is estimated through the two-stage regression as described in Section 2, where feature functions are fixed via their kernel functions. PMMR also models the bridge function using kernel functions, but parameters are learned by moment matching. CEVAE is not a PCL method, however, it represents a state-of-the-art approach in correcting for hidden confounders using observed proxies. The causal graph for CEVAE is shown in Figure 2, and CEVAE uses a VAE Kingma and Welling [2014] to recover the distribution of confounder $U$ from the “proxy” $Q$. We make two modifications to CEVAE to apply it in our setting. First, we include both the treatment-inducing proxy $Z$ and output-inducing proxy $W$ as $Q$ in CEVAE (we emphasize that this does not follow the causal graph in Figure 2, since there exist arrows from $Q$ to $A,Y$). Second, CEVAE is originally used in the setting where $Q$ is conditioned on a particular value, whereas we marginalize $Q$. See Appendix F.4 for the choice of the network structure and hyper-parameters. We tuned the regularizers $\lambda_{1},\lambda_{2}$ as discussed in Appendix A, with the data evenly split for Stage 1 and Stage 2. We varied the dataset size and ran 20 simulations for each setting. Results are summarized in Figure 4.
@@ 5 Conclusion
In future work, it would be interesting to explore different approaches to learning deep models in the PCL problem. One direction would be to adapt the method of moment matching, which has been studied extensively in the instrumental variable literature [Bennett et al., 2019]. Moreover, Liao et al. [2020] recently developed a novel adversarial method to solve a class of functional equations, which would be also a promising approach in the PCL problem. In the RL context, Tennenholtz et al. [2020] shows an interesting connection between PCL learning and off-policy evaluation in a partially observable Markov decision process, and we believe that adapting DFPV to this setting will be of value.
@@ B.1 Existence of bridge function
we can define $h^{*}(a,w)\coloneqq h^{*}_{a}(w).$ The existence of the bridge function corresponds to the results of Deaner [2018, Lemma 1.1.b] and Miao et al. [2018a, Propisition 1]. For simplicity, in constrast to Deaner [2018], we assume that there is no shared quantity between the proxy variables $W$ and $Z$. Following the proofs of the previous works [Deaner, 2018, Appendix C, p. 57] and [Miao et al., 2018a, Appendix 6], we aim to solve the integral equation in (8).
We will make use of the following theorem on the existence of a solution of a linear integral equation [Kress, 1999, Theorem 15.18].
Proposition 5 ([Kress, 1999, Theorem 15.18]).
@@
Assumption 4 is a minimal requisite to apply Proposition 5. The existing works Miao et al. [2018a], Deaner [2018] assume stronger conditions; the operator $E_{a}$ is assumed to be Hilbert-Schmidt, which implies the compactness. Deaner [2018, Assumption A.1] assumes that the joint distribution of $W$ and $Z$ conditioned on $a$ is absolutely continuous with respect to the product measure of $P_{W|A=a}$ and $P_{Z|A=a},$ and its density is square integrable. The density function serves as the integral kernel of the operator $E_{a}$ whose $L^{2}(P_{W|A=a}\otimes P_{Z|A=a})$-norm corresponds to the Hilbert-Schmidt norm of the operator, where we use $\otimes$ to denote the product measure. On the other hand, in the setting of Miao et al. [2018a], all probability distributions have densities with respect to the Lebesgue measure. The operator $E_{a}$ (denoted by $K_{x}$ in Miao et al. [2018a]) is defined by the relevant densities accordingly (see the paragraph after Lemma 2 of Miao et al. [2018a]). The compactness is, as in Deaner [2018], established by the square-integrability of the integral kernel, which implies that the operator $E_{a}$ is Hilbert-Schmidt (see Condition A1 in Miao et al. [2018a]).
It is easy to see that Assumptions 4 and 5 are required for using Proposition 5. The remaining condition to show is that $\mathbb{E}\left[Y|A=a,Z=\cdot\right]$ is in $N(E_{a}^{*})^{\perp}.$ We show that the structural assumption (Assumption 1) and completeness assumption (Assumption 2) imply the required condition. The result below closely follows Deaner [2018] in that we share the same completeness condition.
@@
The difference between the first condition in Assumption 2 and Condition 3 in Miao et al. [2018a] is in the approach to establishing that the conditional expectation $\mathbb{E}[Y|A=a,Z=\cdot]$ belongs to $N(F_{a})^{\perp}.$ More specifically, Condition 3 in Miao et al. [2018a] is equivalent to having $N(F_{a})=\{0\}$ and so any nontrivial $L^{2}(P_{Z|A=a})$-function is in the orthogonal complement. Assumption 2 does not require $N(F_{a})=\{0\}$ but only implies
@@ B.2 Identifiability
The above corollary corresponds to the identifiablity results in obtained in the previous works Miao et al. [2018a], Deaner [2018]. We follow the proof of Theorem 1.1.a in [Deaner, 2018, Appendix B] (See also, Theorem 1 and Appendix 3 of Miao et al. [2018a]).
@@ C.2 Generalization errors for regression
Here, we bound the generalization errors of both stages using Rademacher complexity bounds [Mohri et al., 2012].
@@
[Theorem 3.3 Mohri et al., 2012, with slight modification] Let $\mathcal{S}$ be a measurable space and $\mathcal{H}$ be a family of functions mapping from $\mathcal{S}$ to $[0,M]$. Given fixed dataset $S=(s_{1},s_{2},\dots,s_{n})\in\mathcal{S}^{n}$, the empirical Rademacher complexity is given by
@@
Following the discussion in Xu et al. [2021, Lemma 1], we can show the following lemma that provides a generalization error bound for Stage 1 regression.
@@
The generalization error for Stage 2 can be shown by similar reasoning as in Xu et al. [2021, Lemma 2].
@@
Let Assumption 7 hold and $\kappa_{1},\kappa_{2}=0$. If ^$\hat{\mathfrak{R}}_{S_{1}}(\mathcal{H}_{1})\to 0$ and ^$\hat{\mathfrak{R}}_{S_{2}}(\mathcal{H}_{2})\to 0$ in probability as the dataset size increases, $-^\hat{h}$ converges to ${h^{*}}$ in probability with respect to $\|\cdot\|_{\mathbb{P}\left(A,W\right)}$.
@@ Appendix D DFPV algorithm with observable confounders
In this appendix, we formulate the DFPCL method when observable confounders are present, building on Tchetgen et al. [2020] and Mastouri and Zhu et al. [Mastouri et al., 2021]. Here, we consider the causal graph given in Figure 5. In addition to variables $(A,Y,Z,W)$, we have an observable confounder $X\in\mathcal{X}$. The structural function $f_{\mathrm{struct}}$ we aim to learn is
@@
Assumption 10 (Structural Assumption [Mastouri et al., 2021]).
@@
Assumption 11 (Completeness Assumption [Mastouri et al., 2021]).
@@
Similar to Mastouri and Zhu et al. [Mastouri et al., 2021], we model
@@ E.3 Experiments with Alternative dSprite Experiments
In the proceeding version, we have employed a different setting using dSprite. Here, we use the different structural function given as
@@
where each element of the matrix $B\in\mathbb{R}^{10\times 4096}$ is generated from $Unif(0.0,1.0)$, and the outcome is generated as
$Y=\frac{1}{2}(\mathtt{posY}-0.5)f_{\mathrm{struct}}(A)+\varepsilon,\varepsilon\sim\mathcal{N}(0,0.5).$
We have used the same generation process for variables $(Z,W,A)$.
This benchmark is less challenging compared to the updated structural function presented in the main body since the structural function is less dependent on the position of the image. However, the result is similar to the updated dSprite experiment presented in the main body as shown in Figure 7.
@@ F.1 Demand Design Experiment
    $Y=\frac{1}{12}(\mathtt{posY}-0.5)\frac{(\mathrm{vec}(\bm{B})^{\top}A)^{2}-3000}{500}+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,0.5).$
@@ F.4 Hyper-parameters and network architectures
For DFPV, we optimize the model using Adam [Kingma and Ba, 2015] with learning rate = 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\varepsilon=10^{-8}$. Regularizers $\lambda_{1},\lambda_{2}$ are both set to 0.1 as a result of the tuning procedure described in Appendix A. Network structure is given in Tables 23.  [^4] In ICLR submission, we applied ReLU activation to the final feature output, which causes numerical instability. Please refer to Appendix G for the old results.
@@
Here, $\bm{\psi}_{p(W,Z|L)},\bm{\psi}_{p(A|L)},\bm{\mu}_{p(Y|A,L)}$ are neural networks. We provide the structure of neural nets in Table 4 and 5. Following the orignal work Louizos et al. [2017], we train all neural nets by Adamax [Kingma and Ba, 2015] with a learning rate of 0.01, which was annealed with an exponential decay schedule. We also performed early stopping according to the lower bound on a validation set. To predict structural function, we obtain $q(L)$ by marginalizing $q(L|A,Z,W,Y)$ by observed data $A,Z,W,Y$. We then output $\hat{f}_{\mathrm{struct}}(a)=\mathbb{E}_{L\sim q(L)}\left[\mathbb{E}_{Y\sim p(Y|A=a,L)}\left[Y\right]\right]$.
@@ 4 FC(16, 8)
Table 3: Network structures of DFPV for dSprite experiments. For the input layer, we provide the input variable. For the fully-connected layers (FC), we provide the input and output dimensions. SN denotes Spectral Normalization [Miyato et al., 2018]. BN denotes Batch Normalization.
@@ 5 FC(16, 1)
Table 5: Network structures of CEVAE for dSprite experiment. For the input layer, we provide the input variable. For the fully-connected layers (FC), we provide the input and output dimensions. SN denotes Spectral Normalization [Miyato et al., 2018]. BN denotes Batch Normalization.
@@ 4 FC(128, 256), ReLU
Structure of $\bm{\psi}_{p(A|L)}$ Layer Configuration 1 Input($L$) 2 FC(20, 64), ReLU 3 FC(64, 128), ReLU 4 FC(128, 256), ReLU

### v5

@@ 3.4 DFPV for Policy Evaluation
PCL methods can estimate not only the structural function, but also other causal quantities, such as Conditional Average Treatment Effect (CATE) [29] and Average Treatment Effect on Treated (ATT) [3, 29]. Beyond the causal context, [32] uses a PCL method to conduct the off-policy evaluation in a partially observable Markov decision process with discrete action-state spaces. Here, we present the application of PCL methods to bandit off-policy evaluation with continuous actions, and propose a method to leverage adaptive feature maps in this context.
@@ 4 Experiments
In this section, we report the empirical performance of the DFPV method. First, we present the results of estimating structural functions; we design two experimental settings for low-dimensional treatments and high-dimensional treatments, respectively. Then, we show the result of applying PCL methods to the bandit off-policy evaluation problem with confounding. We include the results for problems considered in prior work in Appendix E. The experiments are implemented using PyTorch [27]. The code is included in the supplemental material. All experiments can be run in a few minutes on Intel(R) Xeon(R) CPU E5-2698 v4 2.20GHz.
@@
Our second structural function estimation experiment considers high-dimensional treatment variables. We test this using the dSprite dataset [19], which is an image dataset described by five latent parameters (shape, scale, rotation, posX and posY). The images are $64\times 64=4096$-dimensional. Based on this, Xu et al. [35] introduced the causal experiment, where the treatment is each figure, and the confounder is posY. Inspired by this, we consider the PCL setting that learns the same structural functions with nonlinear confounding, which is not possible to handle in the instrumental variable setting. Specifically, the structural function $f_{\mathrm{struct}}$ and outcome $Y$ are defined as
$f_{\mathrm{struct}}(a)=\frac{(\mathrm{vec}(\bm{B})^{\top}a)^{2}-3000}{500},\quad Y=12(\mathtt{posY}-0.5)^{2}f_{\mathrm{struct}}(A)+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,0.5),$
@@
We compare the DFPV method to three competing methods, namely KPV [18], PMMR [18], and an autoencoder approach derived from the CEVAE method [17]. In KPV, the bridge function is estimated through the two-stage regression as described in Section 2, where feature functions are fixed via their kernel functions. PMMR also models the bridge function using kernel functions, but parameters are learned by moment matching. CEVAE is not a PCL method, however, it represents a state-of-the-art approach in correcting for hidden confounders using observed proxies. The causal graph for CEVAE is shown in Figure 2, and CEVAE uses a VAE [11] to recover the distribution of confounder $U$ from the “proxy” $Q$. We make two modifications to CEVAE to apply it in our setting. First, we include both the treatment-inducing proxy $Z$ and output-inducing proxy $W$ as $Q$ in CEVAE (we emphasize that this does not follow the causal graph in Figure 2, since there exist arrows from $Q$ to $A,Y$). Second, CEVAE is originally used in the setting where $Q$ is conditioned on a particular value, whereas we marginalize $Q$. See Appendix F.4 for the choice of the network structure and hyper-parameters. We tuned the regularizers $\lambda_{1},\lambda_{2}$ as discussed in Appendix A, with the data evenly split for Stage 1 and Stage 2. We varied the dataset size and ran 20 simulations for each setting. Results are summarized in Figure 4.
@@ 5 Conclusion
In future work, it would be interesting to explore different approaches to learning deep models in the PCL problem. One direction would be to adapt the method of moment matching, which has been studied extensively in the instrumental variable literature [1]. Moreover, Liao et al. [15] recently developed a novel adversarial method to solve a class of functional equations, which would be also a promising approach in the PCL problem. In the RL context, Tennenholtz et al. [32] shows an interesting connection between PCL learning and off-policy evaluation in a partially observable Markov decision process, and we believe that adapting DFPV to this setting will be of value.
@@ B.1 Existence of bridge function
we can define $h^{*}(a,w)\coloneqq h^{*}_{a}(w).$ The existence of the bridge function corresponds to the results of Deaner [3, Lemma 1.1.b] and Miao et al. [20, Propisition 1]. For simplicity, in constrast to Deaner [3], we assume that there is no shared quantity between the proxy variables $W$ and $Z$. Following the proofs of the previous works [3, Appendix C, p. 57] and [20, Appendix 6], we aim to solve the integral equation in (8).
We will make use of the following theorem on the existence of a solution of a linear integral equation [12, Theorem 15.18].
Proposition 5 ([12, Theorem 15.18]).
@@
Assumption 4 is a minimal requisite to apply Proposition 5. The existing works [20, 3] assume stronger conditions; the operator $E_{a}$ is assumed to be Hilbert-Schmidt, which implies the compactness. Deaner [3, Assumption A.1] assumes that the joint distribution of $W$ and $Z$ conditioned on $a$ is absolutely continuous with respect to the product measure of $P_{W|A=a}$ and $P_{Z|A=a},$ and its density is square integrable. The density function serves as the integral kernel of the operator $E_{a}$ whose $L^{2}(P_{W|A=a}\otimes P_{Z|A=a})$-norm corresponds to the Hilbert-Schmidt norm of the operator, where we use $\otimes$ to denote the product measure. On the other hand, in the setting of Miao et al. [20], all probability distributions have densities with respect to the Lebesgue measure. The operator $E_{a}$ (denoted by $K_{x}$ in [20]) is defined by the relevant densities accordingly (see the paragraph after Lemma 2 of [20]). The compactness is, as in [3], established by the square-integrability of the integral kernel, which implies that the operator $E_{a}$ is Hilbert-Schmidt (see Condition A1 in [20]).
It is easy to see that Assumptions 4 and 5 are required for using Proposition 5. The remaining condition to show is that $\mathbb{E}\left[Y|A=a,Z=\cdot\right]$ is in $N(E_{a}^{*})^{\perp}.$ We show that the structural assumption (Assumption 1) and completeness assumption (Assumption 2) imply the required condition. The result below closely follows [3] in that we share the same completeness condition.
@@
The difference between the first condition in Assumption 2 and Condition 3 in [20] is in the approach to establishing that the conditional expectation $\mathbb{E}[Y|A=a,Z=\cdot]$ belongs to $N(F_{a})^{\perp}.$ More specifically, Condition 3 in [20] is equivalent to having $N(F_{a})=\{0\}$ and so any nontrivial $L^{2}(P_{Z|A=a})$-function is in the orthogonal complement. Assumption 2 does not require $N(F_{a})=\{0\}$ but only implies
@@ B.2 Identifiability
The above corollary corresponds to the identifiablity results in obtained in the previous works [20, 3]. We follow the proof of Theorem 1.1.a in [3, Appendix B] (See also, Theorem 1 and Appendix 3 of [20]).
@@ C.2 Generalization errors for regression
Here, we bound the generalization errors of both stages using Rademacher complexity bounds [24].
@@
[Theorem 3.3 24, with slight modification] Let $\mathcal{S}$ be a measurable space and $\mathcal{H}$ be a family of functions mapping from $\mathcal{S}$ to $[0,M]$. Given fixed dataset $S=(s_{1},s_{2},\dots,s_{n})\in\mathcal{S}^{n}$, the empirical Rademacher complexity is given by
@@
Following the discussion in Xu et al. [35, Lemma 1], we can show the following lemma that provides a generalization error bound for Stage 1 regression.
@@
The generalization error for Stage 2 can be shown by similar reasoning as in Xu et al. [35, Lemma 2].
@@
Let Assumption 7 hold and $\kappa_{1},\kappa_{2}=0$. If ^$\hat{\mathfrak{R}}_{S_{1}}(\mathcal{H}_{1})\to 0$ and ^$\hat{\mathfrak{R}}_{S_{2}}(\mathcal{H}_{2})\to 0$ in probability as the dataset size increases, $\hat{h}$ converges to ${h^{*}}$ in probability with respect to $\|\cdot\|_{\mathbb{P}\left(A,W\right)}$.
@@ Appendix D DFPV algorithm with observable confounders
In this appendix, we formulate the DFPCL method when observable confounders are present, building on Tchetgen et al. [31] and Mastouri and Zhu et al. [18]. Here, we consider the causal graph given in Figure 5. In addition to variables $(A,Y,Z,W)$, we have an observable confounder $X\in\mathcal{X}$. The structural function $f_{\mathrm{struct}}$ we aim to learn is
@@
Assumption 10 (Structural Assumption [18]).
@@
Assumption 11 (Completeness Assumption [18]).
@@
Similar to Mastouri and Zhu et al. [18], we model
@@ E.3 Experiments with Alternative dSprite Experiments
In the preceding version of this document, we employed a different experimental dSprite setting. We used the structural function
@@
where each element of the matrix $B\in\mathbb{R}^{10\times 4096}$ was generated from $Unif(0.0,1.0)$, and the outcome was generated as
$Y=12(\mathtt{posY}-0.5)^{2}f_{\mathrm{struct}}(A)+\varepsilon,\varepsilon\sim\mathcal{N}(0,0.5).$
The generation generation process for variables $(Z,W,A)$ is unchanged in the present document.
We now describe a limitation of this setting. [^4] We thank Olawale Salaudeen for alerting us to this issue. Sprite images $A$ have a small number of pixels with value $1$, and many pixels with value 00. Each entry of $BA$ thus represents a sum of uniformly distributed independent random variables from $B$ corresponding to the nonzero entries of $A$. For this reason, the position of $A$ is very difficult to recover from $BA$, since this would require memorizing the specific sum of uniform random values for each sprite position. In practice, the structural function effectively appears as a constant function with additional noise due to $\|BA\|_{2}^{2}$. Nonetheless, even recovery of this (effectively) constant function benefits from the proxy setting. Results of these earlier experiments are show in Figure 7.
@@ F.1 Demand Design Experiment
    $Y=12(\mathtt{posY}-0.5)^{2}\frac{(\mathrm{vec}(\bm{B})^{\top}A)^{2}-3000}{500}+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,0.5).$
@@ F.4 Hyper-parameters and network architectures
For DFPV, we optimize the model using Adam [10] with learning rate = 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\varepsilon=10^{-8}$. Regularizers $\lambda_{1},\lambda_{2}$ are both set to 0.1 as a result of the tuning procedure described in Appendix A. Network structure is given in Tables 23.  [^5] In ICLR submission, we applied ReLU activation to the final feature output, which causes numerical instability. Please refer to Appendix G for the old results.
@@
Here, $\bm{\psi}_{p(W,Z|L)},\bm{\psi}_{p(A|L)},\bm{\mu}_{p(Y|A,L)}$ are neural networks. We provide the structure of neural nets in Table 4 and 5. Following the orignal work [17], we train all neural nets by Adamax [10] with a learning rate of 0.01, which was annealed with an exponential decay schedule. We also performed early stopping according to the lower bound on a validation set. To predict structural function, we obtain $q(L)$ by marginalizing $q(L|A,Z,W,Y)$ by observed data $A,Z,W,Y$. We then output $\hat{f}_{\mathrm{struct}}(a)=\mathbb{E}_{L\sim q(L)}\left[\mathbb{E}_{Y\sim p(Y|A=a,L)}\left[Y\right]\right]$.
@@ 4 FC(16, 8)
Table 3: Network structures of DFPV for dSprite experiments. For the input layer, we provide the input variable. For the fully-connected layers (FC), we provide the input and output dimensions. SN denotes Spectral Normalization [23]. BN denotes Batch Normalization.
@@ 5 FC(16, 1)
Table 5: Network structures of CEVAE for dSprite experiment. For the input layer, we provide the input variable. For the fully-connected layers (FC), we provide the input and output dimensions. SN denotes Spectral Normalization [23]. BN denotes Batch Normalization.
@@ 4 FC(128, 256), ReLU
Structure of $\bm{\psi}_{p(A|L)}$
Layer Configuration
1 Input($L$)
2 FC(20, 64), ReLU
3 FC(64, 128), ReLU
4 FC(128, 256), ReLU